HUBBLE SPACE TELESCOPE DAILY REPORT #2575 PERIOD COVERED: 0000Z (UTC) 03/07/00 - 0000Z (UTC) 03/08/00 Daily Status Report as of 068/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 8216 (T Tauri Star Snapshot Survey II. Completing the CTTs Sample) The WF/PC-2 was used to make observations of the T-Tauri class star known as WSB-28 as part of Snapshot Survey II. This database will be of fundamental importance to understanding the nature of planet-forming environments and their frequency in the galaxy. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed STIS/CCD 8439 (Bias Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no reported problems. 1.3 Completed Three Sets of WF/PC-2 8460 (Cycle 8 Supplemental Darks pt2/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations completed with no reported problems. 1.4 Completed Two Sets of STIS/CCD 8437 (Dark Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. The proposal completed nominally. 1.5 Completed WF/PC-2 8230 (Snapshot Survey of Dynamically Close Galaxy Pairs from z=0.1 to z=0.5) The WF/PC-2 was used to make snapshot survey observations of galactic cluster Gal-Clus-0. This is to obtain snapshot images of a large, well-defined, and dynamically selected sample of close galaxy pairs at moderate redshift. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Six Sets of WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed STIS/CCD 8299 (Mapping the Dynamics of the Quasar 3C 48) The Space Telescope Imaging Spectrograph (CCD) was used to determine the relation between the properties of the radio source and those of the {optical} line-emitting gas in quasars and radio galaxies. The observations completed with no reported problems. 1.09 Completed STIS/CCD 8796 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation IV) The Space Telescope Imaging Spectrograph (CCD) was used in the parallel mode to make some low galactic latitude archive observations. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed STIS/CCD 8240 (Wilson Masses and Spins of Black Holes in Seyfert Galaxies) The Space Telescope Imaging Spectrograph (CCD) was used to observe about 20 bright Seyfert 1 galaxies that are now known to have very broad {up to 100, 000 km s^-1} Fe KAlpha emission lines near 6.4 keV. The observations completed nominally. 1.11 Completed STIS/CCD 8411 (STIS CCD Spectroscopic Flats C-8) The Space Telescope Imaging Spectrograph (CCD) was used to obtain CCD flats in the spectroscopic mode. The observations were completed as scheduled, and no anomalies were reported. 1.12 Completed STIS/CCD 8135 (Spectroscopy of Gravitational Lens Candidates from the HST Survey of BL Lac Objects) The Space Telescope Imaging Spectrograph (CCD) was used to make spectroscopic observations of two new gravitational lens candidate, 0502+675, discovered in our HST R-band snapshot survey of ~100 BL Lac objects. It contains a double point source with separation of only 0.3 arcsec. A candidate for the lensing galaxy has also found. STIS long-slit spectra of each point source will determine whether these candidates are lenses and also measure the redshift of the possible lensing galaxy. The proposal completed nominally. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 5 Successful: 5 Scheduled Re-acquisitions: 9 Successful: 9 2.2 FHST Updates: Scheduled: 15 Successful: 15 2.3 Operations Notes: Using ROP SR-1A, the SSR EDAC error counter was cleared one time. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW